4. Calculate
a. q when a system does 54J of work and its energy decreases by 72 J
b. ∆E for a gas that releases 38J of heat and has 102J of work done on it.
1
Expert's answer
2020-10-30T06:43:20-0400

a) The first law of thermodynamics helps us in calculating the amount of heat absorbed or released.

From the formula;"\Delta_E" "=q+w"

"q=\Delta_E" "-w"

"\therefore" "q=(-72J)-(-54J)"

"=-18J"

b) Using the first law of thermodynamics helps us to calculate the change in internal energy. "q" in this case is a negative value since the system releases energy.

"\Delta_E=q+w"

"=(-38J)+(+102J)=+64J"